# Digital Over-the-Air Federated Learning in Multi-Antenna Systems

### v2

@@ II-A Digital pre-processing at the devices
where $\hat{\bm{w}}_{k,t}\in{\bf R}^{W}$ is a modulated symbol vector with $W$ being the number of symbols, and $l\left(\cdot\right)$ denotes the digital pre-processing function that combines decimal-to-binary conversion and digital modulation where the decimal-to-binary conversion is used to represent each numerical FL parameter with a binary coded bit-interleaved vector and the digital modulation is used to modulate several binary bits as a symbol [43]. For convenience, the modulated signal $\hat{\bm{w}}_{k,t}$ is normalized (i.e., $|\hat{\bm{w}}_{k,t}|=1$). We use rectangular $M$-quadrature-amplitude modulation (QAM) for digital modulation and it can be extended to other types of digital modulation schemes.
@@ III Solution for Problem (9)
III-A Analysis of the Convergence of the Designed FL
@@
III-B Prediction of the Local FL Models
@@
Next, we explain the use of neural networks to predict the local FL model updates of all devices. The prediction of FL model parameters at each FL training iteration is considered a regression task since the values of the FL model are continuous. Therefore, we propose to use multilayer perceptrons (MLPs) instead of other neural networks (such as recurrent neural networks) since MLPs are particularly skilled at handling regression tasks [51]. Next, we first explain the components of the proposed ANN-based algorithm. Then, the details to implement this algorithm for predicting each local FL model update are presented.
@@
The proposed MLP-based prediction algorithm consists of three components: a) input, b) a single hidden layer, and c) output, which are defined as follows:
  • Input: The input of the MLP that is implemented by the PS for predicting device $k$’s local FL model is a vector ${\bm{g}}_{t-1}$. As we mentioned in (6), all devices are able to connect with the PS so as to provide the input information for the MLP to predict the local FL models for next iteration.
  • Output: The output of the MLP is a vector ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}$ that represents device $k$’s local FL model update at current iteration $t$. Based on the predicted ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}$, the PS can adjust the transmit and receive beamforming matrices proactively to minimize problem (15).
@@
  • A Single Hidden Layer: The hidden layer of a MLP is used to learn the nonlinear relationships between input vector ${\bm{g}}_{t-1}$ and the output vector ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}$. The weight matrix that represents the connection strength between the input vector and the neurons in the hidden layer is ${\bm{v}}^{\rm in}\in{\bf C}^{D\times V}$ where $D$ is the number of neurons in the single hidden layer. Meanwhile, the weight matrix that captures the strengths of the connections between the neurons in the hidden layer and the output vector is ${\bm{v}}^{\rm out}\in{\bf C}^{V\times D}$.
Having the components of the MLP, next, we introduce the use of the MLP to predict each device’s local FL model update. The states of the neurons in the hidden layer are given by
${\bm{v}}=\sigma\left({\bm{v}}^{\rm in}{\bm{g}}_{k,t-1}+{\bm{b}_{v}}\right),$ (16)
@@
where $\sigma\left(x\right)=\frac{2}{1+{\rm exp}\left(-2x\right)}-1$ and ${\bm{b}_{v}}\in{\bf C}^{D\times 1}$ is the bias. Then, the output of the MLP can be given by
${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}={\bm{v}}^{\rm out}\bm{v}+{\bm{b}_{o}},$ (17)
where ${\bm{b}_{o}}\in{\bf C}^{V\times 1}$ is a vector of bias.
@@
To predict each device’s local FL model update, the MLP must be trained by an online gradient descent method. However, in the considered model, the PS can only obtain $\bm{g}_{t}$ that is directly demodulated from the received signal from all devices. Hence, the PS and the devices must exchange information to train the MLP cooperatively. In particular, at each iteration, device $k$ first generates $\bm{w}_{k,t}$ using its local dataset and $\bm{g}_{t-1}$ received from the PS. Then, device $k$ calculates the training loss of MLP and transmits it to the PS. Based on the value of the training loss, the PS and device $k$ can update its MLP synchronously. Since each device only needs to transmit its training loss, the cost for information exchange can be ignored compared with the transmission of DNN model parameters in FL training.
@@ III-D Implementation and Complexity
Next, we introduce the implementation and complexity of the designed FL algorithm. With regards to the implementation of the proposed algorithm, the PS must a) use MLPs to predict the devices’ local FL models, and b) design the transmit and receive beamforming matrices based on the predicted models. To train the MLPs that are used for the predictions of devices’ local FL models, the PS will use the global FL model $\bm{g}_{t-1}$ that is directly reconstructed from the received symbol vector ${\bm{\hat{s}}_{t-1}}$ at iteration $t-1$. Since ${\bm{\hat{s}}_{t-1}}$ that is transmitted from all devices via AirComp is originally used for the update of the global FL model, the PS does not require any additional information for training MLPs. To design the optimal transmit and receive beamforming matrices, the PS requires the maximal transmit power $P_{0}$ and the MIMO channel vector $\bm{H}_{k}$ of each device $k$. Since $P_{0}$ is a fixed scalar, the data size of which is much smaller than the data size of the local FL models that the devices must transmit to the PS during each iteration. In consequence, we can ignore the overhead of each device transmitting $P_{0}$ to the PS. With regards to $\bm{H}_{k}$, the PS can use channel estimation methods to learn $\bm{H}_{k}$ over each uplink channel so as to design optimal transmit and receive beamforming matrices.
@@
To satisfy the synchronization requirement in the proposed AirComp framework, it is necessary to incorporate clock synchronization for precise control of information transmission timing and channel state information (CSI) feedback for accurate estimation of transmission delay. For clock synchronization, the PS can broadcast a shared block to all devices to achieve time calibration, as down in previous studies [53, 54, 55]. For accurate estimation of transmission delay, the PS can obtain the CSI directly from the uplink pilot signals transmitted from devices, as done in previous studies [56, 57, 58].
We identify two components of our algorithm that could potentially impact complexity and latency: (1) the MLP and (2) optimizing $\bm{A}_{t}$ and $\bm{B}_{t}$. The overhead associated with the MLPs is small for the following reasons. First, the MLP is trained via an offline method, and it only needs to be trained once. Second, the MLP is trained at the PS, which we can reasonably assume has sufficient computational resources [59]. Third, the training complexity of the MLP can be negligible compared to FL training. Specifically, the computational complexity of training a MLP that consists of only three fully connected layers lies in the size of input $\bm{g}_{t-1}$ and output ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}$, as well as the number of the neurons in the hidden layer. The sizes of the global FL model $\bm{g}_{t-1}$ and each local FL model ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}$ are both $V$, as defined in (1), while the number of neurons in the hidden layer is $D$, as defined in (16). Hence, the computational complexity is ${\mathcal{O}}(2\times V\times D)$, which implies that the training complexity of the MLP can be negligible compared to FL training. The relatively low-training complexity of the MLP also implies that the additional computational latency introduced by MLP inference is relatively minor. Communication of the model parameters between the clients and the PS is much more time consuming than forward passing the signal through an ANN at the PS.
@@
Hence, we only analyze the complexity of a dual method that is used to optimize $\bm{A}_{t}$ and $\bm{B}_{t}$ which has to be done for each iteration and thus the complexity scales lies in the number of iterations required to converge. For finding optimal $\bm{A}_{t}$ and $\bm{B}_{t}$, problem (22) can be solved by a traditional augmented Lagrangian method that approaches the optimal solution via alternating updating $\bm{A}_{t}$, $\bm{B}_{t}$, and the Lagrangian multiplier vector. Obviously, the introduced Lagrangian multiplier vector consists of $K$ constraints where $K$ is the number of devices in the considered FL framework. Hence, the PS is required to sequentially update $K$ Lagrangian multipliers, $\bm{A}_{t}=[\bm{A}_{1,t},\cdots,\bm{A}_{K,t}]$, and $\bm{B}_{t}$ at each iteration. Let $L_{O}$ be the number of iterations until the traditional augmented Lagrangian method converges, the complexity is $\mathcal{O}(L_{O}K^{2})$.
@@
IV Simulation Results and Analysis
We consider a circular network area with a radius $r=1500$ m with one PS at its center serving $K=20$ uniformly distributed devices. In particular, the PS allocates 14 subcarriers to all devices and the bandwidth of each subcarrier is 15 kHz. By default, the channels between the PS and devices are modeled as independent and identically distributed with Rayleigh fading and a pathloss exponent of $\beta=2$. The other default settings of parameters used in simulations are listed in Table II.
@@ IV Simulation Results and Analysis
To evaluate the performance of the proposed FL, the MNIST dataset [60], Fashion-MNIST dataset [61], and CIFAR-10 dataset [62] are used as ML tasks. In particular, for MNIST, we adopt a fully-connected neural network (FNN) that consists of two full-connection layers with 7840 (=28$\times$28$\times$10) model parameters. For Fashion-MNIST, each FL model consists of two convolutional layers and three fully-connected layers with 43500 model parameters. For CIFAR-10, each local FL model is a standard ResNet-18 that consists of 16 convolutional layers and two fully-connected layers with $1.17\times 10^{7}$ model parameters. The PS uses one MLP that consists of three layers to predict the FL gradient vector of each device.
Throughout the simulations, we will consider data samples to be allocated across devices in either an independent and identically distributed (IID) or non-IID manner. In the IID data setting, each device is allocated datapoints from all available class labels, while in the non-IID case, they each only receive datapoints from a fraction of the labels [63, 64].
It is worth noting that the convolution kernels exhibit sensitivity to errors, which significantly degrade the performance of the FL model after quantization, as demonstrated in the simulations. However, considering the relatively small number of parameters in convolution kernels, it is viable to utilize traditional methods such as orthogonal frequency division multiple access for transmitting the convolution kernel parameters in full precision.
In our experiments, all FL algorithms are considered to have converged when the value of the FL loss variance calculated over five consecutive iterations is less than 0.001.
@@
$r$ 1500 m $D$ 128 $L_{0}$ 100
IV-A Training Speed and Accuracy Comparisons
@@
We first conduct several experiments comparing our methodology to the baselines in terms of convergence speed and accuracy. In Fig. 3, we show how the test accuracy of all considered algorithms changes over communication rounds, on the Fashion-MNIST task, for the IID data setting. In this figure, we can see that, the proposed algorithm improves the test accuracy by up to 10.5% and 27.5% compared to analog FL and BPSK FL, respectively. From Fig. 3, we can also see that the performance of analog FL experiences noticeable fluctuations. This is due to the noise over wireless channels introducing dynamic errors into FL parameter transmission process, thus affecting FL test accuracy. We also see that the proposed method without using MLP for predicting FL gradients cannot converge, which verifies that the optimal beamforming matrices design depends on the prediction of FL gradients.
Figure 4: Test accuracy of the proposed AirComp-based FL system over time on the Fashion-MNIST task, for an IID data allocation. The overall trends are similar to those observed in Fig. 3.
Fig. 4 shows how the test accuracy changes over time elapsed, measured in seconds, as opposed to iterations. Here, the elapsed time consists of the FL model training time, inference time that is used for predicting FL parameters, and the model transmission time. The overall trends are similar to those observed in Fig. 3: compared to BPSK FL, while our proposed methodology converges more slowly, it ends up improving the test accuracy by up to 27%. This is because the proposed FL uses more bits instead of one bit in BPSK FL to represent each FL parameter, thus increasing the dynamics of global FL model generation. In addition, this figure also shows that the proposed algorithm can reduce the convergence time by 16% compared to analog FL. This finding underscores the fact that although the utilization of MLP introduces additional inference latency, it ultimately results in a decrease in the overall convergence time of FL. We will further investigate the latency vs. accuracy tradeoff in Tables III and IV.
@@
In Fig. 6, we repeat the experiment from Fig. 5, but this time under a non-IID setting, where the data samples of each device are drawn from 6 (instead of 10) classes. From this figure, we see that the proposed FL with 64-QAM can improve the test accuracy by up to 20% and 10% compared with analog FL and ChannelComp, similar to in the IID case. Fig. 6 also shows that under noiseless channels, the accuracy of the AirComp-based system is 8% better, consistent with the IID case.
@@
Finally, in Tables III and IV, we compare the performance of the proposed method with that of analog FL and BPSK FL in terms of (i) latency per iteration, (ii) overall training time, and (iii) converged test accuracy, on the MNIST and CIFAR-10 datasets, respectively, for the non-IID cases. Overall, compared to the baselines, we see that while our method incurs a higher latency in each round, it has a smaller latency across the entire training process compared with analog FL, since it requires less total rounds to converge. Employing the MLP within our method introduces additional latency and computational overhead, but aids in designing the beamforming matrices to reduce transmission errors, thereby reducing the total number of rounds required. BPSK FL has the lowest latencies of all, but suffers from poor testing accuracy at convergence, due to the fact that it quantizes the FL parameters to a single bit. We observe too that the per-round and total latencies are higher in CIFAR-10, which is the most complex of the tasks, with the largest ML model being trained and transmitted.
TABLE III: Latency vs. accuracy tradeoff for training on the MNIST dataset.
  Averaged Latency for Each Round Averaged Latency for Entire Training Converged Test Accuracy Proposed 0.0231 s 0.18 s 91% Analog FL 0.0185 s 0.37 s 88% BPSK FL 0.0135 s 0.07 s 85%
@@
  Averaged Latency for Each Round Averaged Latency for Entire Training Converged Test Accuracy Proposed 0.394 s 15.76 s 65% Analog FL 0.298 s 17.88 s 50% BPSK FL 0.276 s 11.04 s 26%
IV-B Robustness of MLP Predictor

### v3

@@ II-A Digital Pre-Processing at the Devices
where $\hat{\bm{w}}_{k,t}\in{\bf R}^{W}$ is a modulated symbol vector with $W$ being the number of symbols, and $l\left(\cdot\right)$ denotes the digital pre-processing function that combines decimal-to-binary conversion and digital modulation, where the decimal-to-binary conversion is used to represent each numerical FL parameter with a binary coded bit-interleaved vector, and the digital modulation is used to modulate several binary bits as a symbol [43]. For convenience, the modulated signal $\hat{\bm{w}}_{k,t}$ is normalized (i.e., $|\hat{\bm{w}}_{k,t}|=1$). We use rectangular $M$-quadrature-amplitude modulation (QAM) for digital modulation and it can be extended to other types of digital modulation schemes.
@@ III Solution for Problem (9)
III-A Convergence Analysis of Designed FL
@@
III-B Prediction of Local FL Models
@@
Next, we explain the use of neural networks to predict the local FL model updates of all devices at the PS. Formally, the task is to predict parameters of device $k$’s FL model update at time $t$, i.e., ${\rm\Delta}{\bm{w}}_{k,t}\in{\bf R}^{V\times 1}$, from the global aggregation available at the PS at time $t-1$, i.e., ${\bm{g}}_{t-1}\in{\bf R}^{V\times 1}$. This is considered a regression task since the values of the FL model are continuous. Therefore, we propose to use multilayer perceptrons (MLPs), a standard type of artificial neural network (ANN) for handling regression tasks [51].
@@
Our proposed MLP-based prediction algorithm consists of three layers: (a) input layer, (b) a single hidden layer, and (c) output layer. These components are defined as follows:
  • Input layer: The input to the MLP is a vector ${\bm{g}}^{\prime}_{t-1}\in{\bf R}^{V^{\prime}\times 1}$ that represents the previous aggregated results of the FL parameters being predicted. This is a subset of ${\bm{g}}_{t-1}\in{\bf R}^{V\times 1}$, where the number of parameters $V^{\prime}\leq V$ is adjustable. Since the total parameter dimension $V$ may be large (e.g., for the CIFAR-10 experiments in Sec. IV, $V=1.17\times 10^{7}$), selecting a smaller number $V^{\prime}$ to predict can lead to complexity reduction advantages. As we mentioned in (6), all devices are able to connect with the PS so as to provide the input information for the MLP to predict the local FL models for next iteration.
  • Output layer: The output of the MLP is a vector ${\rm\Delta}{\widetilde{\bm{w}}^{\prime}}_{k,t}\in{\bf R}^{V^{\prime}\times 1}$ that represents prediction of the $V^{\prime}$ parameters in device $k$’s local FL model update for the current iteration $t$. The resulting estimate ${\rm\Delta}{\widetilde{\bm{w}}}_{k,t}\in{\bf R}^{V\times 1}$ of ${\rm\Delta}{\bm{w}}_{k,t}$ used in the beamforming design in Section III-C is then a combination of (a) the $V^{\prime}$ parameter predictions ${\rm\Delta}{\widetilde{\bm{w}}^{\prime}}_{k,t}$ that the ANN has made, and (b) the other $V-V^{\prime}$ parameters in ${\bm{g}}_{t-1}$ that are not part of the ANN.
@@
  • Single hidden layer: We employ a single hidden layer of dimension $D$ to learn the nonlinear relationships between the input ${\bm{g}}^{\prime}_{t-1}$ and the output ${\rm\Delta}{\widetilde{\bm{w}}}^{\prime}_{k,t}$. The weight matrix between the input vector and the neurons in the hidden layer for device $k$ is denoted ${\bm{v}}^{\rm in}_{k}\in{\bf R}^{V^{\prime}\times D}$. Meanwhile, the weight matrix between the neurons in the hidden layer and the output vector is denoted ${\bm{v}}^{\rm out}_{k}\in{\bf R}^{D\times V^{\prime}}$.
Based on this, the states of the neurons in the hidden layer are given by
${\bm{v}}_{k,t}=\sigma\left({\bm{v}}_{k}^{\rm in}{\bm{g}}^{\prime}_{t-1}+{\bm{b}_{k}^{v}}\right),$ (16)
@@
where $\sigma\left(x\right)=\frac{2}{1+{\rm exp}\left(-2x\right)}-1$ is a sigmoidal activation and ${\bm{b}_{k}^{v}}\in{\bf R}^{D^{\prime}\times 1}$ is the bias vector. Then, the output of the MLP is given by
${\rm\Delta}{\widetilde{\bm{w}}^{\prime}}_{k,t}={\bm{v}}_{k}^{\rm out}\bm{v}_{k,t}+{\bm{b}_{k}^{o}},$ (17)
where ${\bm{b}_{k}^{o}}\in{\bf R}^{V^{\prime}\times 1}$ is another bias vector.
@@
The MLP is trained through an online gradient descent method in parallel with the FL model updating. However, in the considered model, the PS only has access to the value of $\bm{g}_{t}$ that is directly demodulated from the received signal from all devices. Hence, the PS and the devices must exchange information to train the MLP for each device cooperatively. In particular, at each iteration, device $k$ first calculates its local update ${\rm\Delta}\bm{w}_{k,t}$ and $\bm{g}_{t-1}$ received from the PS, and extracts ${\rm\Delta}\bm{w}_{k,t}^{\prime}\in{\bf R}^{V^{\prime}\times 1}$ from ${\rm\Delta}\bm{w}_{k,t}$. Then, it calculates the gradient of the MLP parameters with respect to the MLP’s prediction of ${\rm\Delta}\bm{w}_{k,t}^{\prime}$ from $\bm{g}^{\prime}_{t-1}$. Finally, device $k$ transmits these gradients to the PS, which updates the MLP parameters accordingly. We will see in Sec. IV that the per-iteration latency introduced by this MLP procedure is overcome by the reduction in the number of training rounds needed for convergence, so long as the size of the MLP is limited.
@@ III-D Implementation and Complexity
Next, we discuss the implementation and complexity of the designed FL algorithm. With regards to the implementation of the proposed algorithm, the PS must a) use MLPs to predict the devices’ local FL model update, and b) design the transmit and receive beamforming matrices based on the predicted updates. To train the MLPs that are used for the predictions of devices’ local FL model update, the PS will use the global FL model $\bm{g}_{t-1}$ that is directly reconstructed from the received symbol vector ${\bm{\hat{s}}_{t-1}}$ at iteration $t-1$. Additionally, the PS needs the gradients of the MLP parameters ${\bm{v}}_{k}^{\rm in},{\bm{v}}_{k}^{\rm out}$ calculated on the device. To design the optimal transmit and receive beamforming matrices, the PS requires the maximal transmit power $P_{0}$ and the MIMO channel vector $\bm{H}_{k}$ of each device $k$. We ignore the overhead of each device transmitting $P_{0}$ to the PS since it is a scalar. With regards to $\bm{H}_{k}$, the PS can use channel estimation methods to learn $\bm{H}_{k}$ over each uplink channel so as to design optimal transmit and receive beamforming matrices.
@@
To satisfy the synchronization requirement in the proposed AirComp framework, it is necessary to incorporate clock synchronization for precise control of information transmission timing and channel state information (CSI) feedback, for accurate estimation of transmission delay. For clock synchronization, the PS can broadcast a shared block to all devices to achieve time calibration, as in previous studies [53, 54, 55]. For accurate estimation of transmission delay, the PS can obtain the CSI directly from the uplink pilot signals transmitted from devices, as done in previous studies [56, 57, 58].
We identify two components of our algorithm that could potentially impact complexity and latency: (1) the MLP and (2) optimizing $\bm{A}_{t}$ and $\bm{B}_{t}$. The training complexity of the MLP can be made small compared to the complexity of FL training. Specifically, the computational complexity of the MLP lies in the size of input $\bm{g}^{\prime}_{t-1}$ and output ${\rm\Delta}{\widetilde{\bm{w}}}^{\prime}_{k,t}$, as well as the number of the neurons in the hidden layer. As discussed in Sec. III-B, the sizes of $\bm{g}^{\prime}_{t-1}$ and ${\rm\Delta}{\widetilde{\bm{w}}}^{\prime}_{k,t}$ are both $V^{\prime}$, while the number of neurons in the hidden layer is $D$. Hence, the computational complexity is ${\mathcal{O}}(2V^{\prime}D)$, which implies that the training complexity of the MLP can be controlled directly based on limiting the size of $V^{\prime}$ and $D$. In Sec. IV, for the ML task, we will set MLP to predict parameters for the last layer of the neural network, so that $V^{\prime}\ll V$. We can further select $D$ such that $2V^{\prime}D\ll V$, i.e., so the MLP is substantially smaller than the FL model.
@@
Regarding the optimization of $\bm{A}_{t}$ and $\bm{B}_{t}$, the complexity scales in the number of iterations required for the solver to converge. For finding optimal $\bm{A}_{t}$ and $\bm{B}_{t}$, problem (22) can be solved by a traditional augmented Lagrangian method that approaches the optimal solution via alternating updating $\bm{A}_{t}$, $\bm{B}_{t}$, and the Lagrangian multiplier vector. The introduced Lagrangian multiplier vector consists of $K$ constraints, where $K$ is the number of devices in the considered FL framework. Hence, the PS is required to sequentially update $K$ Lagrangian multipliers, $\bm{A}_{t}=[\bm{A}_{1,t},\cdots,\bm{A}_{K,t}]$, and $\bm{B}_{t}$ at each iteration. Letting $L_{O}$ be the number of iterations until the augmented Lagrangian method converges, the complexity is $\mathcal{O}(L_{O}K^{2})$.
@@
IV Experimental Results and Discussion
IV-A Simulation Setup
We consider a circular network area with a radius $r=1500$ m with one PS at its center serving $K=20$ uniformly distributed devices. In particular, the PS allocates 64 subcarriers to all devices and the bandwidth of each subcarrier is 15 kHz. By default, the channels between the PS and devices are modeled as independent and identically distributed with Rayleigh fading and a pathloss exponent of $\beta=2$. The other default settings of parameters used in simulations are listed in Table II.
@@ IV-A Simulation Setup
The Fashion-MNIST dataset [59] and CIFAR-10 dataset [60] are used as ML tasks in our performance evaluation. For Fashion-MNIST, each local FL model consists of five convolutional layers and one fully-connected layer, with $V=9.2\times 10^{5}$ total model parameters. For CIFAR-10, each local FL model is a standard ResNet-18 that consists of 17 convolutional layers and one fully-connected layer, with $V=1.17\times 10^{7}$ total model parameters.
For the MLP described in Sec. III-B, we focus on predicting the parameters of the last layer in each ML model (i.e., the fully connected layers). This leads to $V^{\prime}=640$ for both the CNN in Fashion-MNIST and the Resnet-18 in CIFAR-10. To restrict the ANN’s total size to a small fraction of the ML model in each case, we set $D=160$ for Fashion-MNIST and $D=320$ for CIFAR-10, in line with CIFAR-10’s higher complexity (leading to the MLP being 22% and 3% of the model sizes, respectively). As we will see in Sec. IV-C, we find that this retains a strong MSE performance for each dataset. Similar considerations can be made for other ML tasks.
Throughout the simulations, we will consider data samples to be allocated across devices in either an independent and identically distributed (IID) or non-IID manner. In the IID data setting, each device is allocated datapoints from all 10 class labels, while in the non-IID case, they each only receive datapoints from a fraction of the labels, assumed to be 6 by default [61, 62].
It is worth noting that the convolution kernels exhibit sensitivity to errors, which significantly degrade the performance of the FL model, as will be demonstrated in the simulations. However, considering the relatively small number of parameters in convolution kernels, it is viable to utilize traditional methods such as orthogonal frequency division multiple access for transmitting the kernel parameters.
Finally, in some of our experiments (Fig. 4, Table III, Table IV, Fig. 9), we will conduct accuracy comparisons over time/latency incurred rather than communication rounds. These calculations employ standard models for communication and computation delays [63, 64], which we present in Appendix B. For experiments that refer to convergence (in particular, Tables IIIIV), algorithms are considered to have converged when the value of the FL loss variance calculated over five consecutive iterations is less than $0.001$.
@@
$r$ 1500 m $D$ 160 $L_{0}$ 100
IV-B Training Speed and Accuracy Comparisons
@@
We first conduct several experiments comparing our methodology to the baselines in terms of convergence speed and accuracy. In Fig. 3, we show how the test accuracy of all considered algorithms changes over communication rounds, on the Fashion-MNIST task, for the IID data setting. In this figure, we can see that, the proposed algorithm improves the test accuracy by up to 15.5% and 24.5% compared to analog FL and BPSK FL, respectively. From Fig. 3, we can also see that the performance of analog FL experiences noticeable fluctuations. This is due to the noise over wireless channels introducing dynamic errors into FL parameter transmission process, thus affecting FL test accuracy. We also see that the proposed method without using MLP for predicting FL gradients cannot converge, which verifies that the optimal beamforming matrices design depends on the prediction of FL gradients.
Figure 4: Test accuracy of the proposed AirComp-based FL system over time on the Fashion-MNIST task, for an non-IID data allocation. The overall trends are similar to those observed in Fig. 3.
Fig. 4 shows how the test accuracy changes over time elapsed, measured in seconds, as opposed to iterations. Here, the elapsed time consists of the FL model training time, inference time that is used for predicting FL parameters, and the model transmission time. The overall trends are similar to those observed in Fig. 3: compared to BPSK FL, while our proposed methodology converges more slowly, it ends up improving the test accuracy by up to 27%. This is because the proposed FL uses more bits instead of one bit in BPSK FL to represent each FL parameter, thus increasing the dynamics of global FL model generation. In addition, this figure also shows that the proposed algorithm can reduce the convergence time by 20.3% compared to analog FL. This finding underscores the fact that although the utilization of MLP introduces additional inference latency, it ultimately results in a decrease in the overall convergence time of FL. We will further investigate the latency vs. accuracy tradeoff in Tables III and IV.
@@
In Fig. 6, we repeat the experiment from Fig. 5, but this time under a non-IID setting, where the data samples of each device are drawn from 6 (instead of 10) classes. From this figure, we see that the proposed FL with 64-QAM can improve the test accuracy by up to 18% and 13% compared with analog FL and ChannelComp, similar to in the IID case. Fig. 6 also shows that under noiseless channels, the accuracy of the AirComp-based system is 6% better, consistent with the IID case.
@@
Finally, in Tables III and IV, we compare the performance of the proposed method with that of analog FL and BPSK FL in terms of (i) latency per iteration, (ii) overall training time, (iii) average rounds to convergence, and (iv) converged test accuracy, on the Fashion-MNIST and CIFAR-10 datasets, respectively, for the non-IID cases. Compared to analog FL we see that while our method incurs a higher latency in each round (e.g., $1.26$ vs. $1.03$ sec in Fashion-MNIST), it has a smaller latency across the entire training process (e.g., $18.90$ vs. $23.69$ sec in Fashion-MNIST), since it requires less total rounds to converge. Employing the MLP within our method introduces additional latency and computational overhead, but aids in designing the beamforming matrices to reduce transmission errors, thereby reducing the total number of rounds required. BPSK FL has the lowest latencies of all, but suffers from poor testing accuracy at convergence, due to the fact that it quantizes the FL parameters to a single bit. We observe too that the per-round and total latencies are higher in CIFAR-10, which is the most complex of the tasks, with the largest ML model being trained and transmitted.
TABLE III: Latency vs. Accuracy Tradeoff for Training on the Fashion-MNIST Dataset.
  Avg. Lat. for Each Round Avg. Lat. for Entire Training Avg. Rounds to Converge Converged Test Acc. Proposed 1.26 s 18.90 s 15 90% Analog FL 1.03 s 23.69 s 23 76% BPSK FL 0.21 s 1.05 s 5 67%
@@
  Avg. Lat. for Each Round Avg. Lat. for Entire Training Avg. Rounds to Converge Converged Test Acc. Proposed 21.46 s 966 s 45 73% Analog FL 20.92 s 1255 s 60 52% BPSK FL 2.95 s 118 s 40 31%
IV-C Robustness of MLP Predictor